OpenStudy (anonymous):

A truck on a straight road starts from rest, accelerating at 2.00 m/s2 until it reaches a speed of 45.0 m/s. Then the truck travels for 67.0 s at constant speed until the brakes are applied, stopping the truck in a uniform manner in an additional 5.00 s. A) How long is the truck in motion? (already solve-94.5 sec) B) What is the average velocity of the truck for the motion described?

OpenStudy (cwrw238):

for the first part of the journey use the constant acceleration equation v^2 = u^2 + 2as where u= 0, v = 45 m/s , a = 2 and s = distance travelled distance for middle part = 67 * 45 m for the last part initial speed u = 45 m/s, v = 0, t = 5 to find distance s use the formula s =(u + v)t/2 add these 3 distances and divide by 94.5 secs to find average velocity